The chief accountant of Mobendix Corporation is studying how lot size and manhours are related. She wants to develop a prediction equation showing how varying lot sizes impact the manhours she should anticipate. Provide all necessary assistance. The data obtained recently is shown below Row Y (MHRS) 73 50 128 170 87 108 135 69 148 132 X (LTSZ) 30 20 60 80 4 0 50 60 30 70 60 2 6 9 10
Answer #1

Y- M (X-Mx)2 X- Mx -20 - 30 18ク 30 -10 -37 -60 18 400 900 100 900 100 740 1800 180 1800 230 -23 10 -20 20 10 25 -41 38 100 400 400 100 250 820 760 220 SS: 3400 SP: 6800

Sum of X = 500
Sum of Y = 1100
Mean X = 50
Mean Y = 110
Sum of squares (SSX) = 3400
Sum of products (SP) = 6800

Regression Equation = ŷ = bX + a

b = SP/SSX = 6800/3400 = 2

a = MY - bMX = 110 - (2*50) = 10

ŷ = 2X + 10

250 200 150 100 50母 20 60 80 100 40 x-Independent Regression Line (y = 2x + 10)